OpenStudy (anonymous):

1/7x=1/9x extraneous solution

hero (hero):

What about it?

OpenStudy (anonymous):

it said to solve the equation and check for extraneous solutions

hero (hero):

I assume you are given \[\frac{1}{7x} = \frac{1}{9x}\] and not \[\frac{1}{7}x = \frac{1}{9}x\]

OpenStudy (anonymous):

yes the first one is what i have @Hero

hero (hero):

Do you know how to cross multiply?

OpenStudy (anonymous):

yes

hero (hero):

What did you get after cross multiplication?